OpenStudy (baewolfstar):

Find f(5) for this sequence: f(1) = 2 and f(2) = 5, f(n) = f(1) + f(2) + f(n - 1), for n > 2. f(5) = ______ Numerical Answers Expected! Answer for Blank 1: Please help me i forgot how to do this lolol

OpenStudy (phi):

they tell you f(n)= f(1) + f(2) + f(n-1) f(1) + f(2) will always be 2+5= 7, so the rule can be written f(n)= f(n-1) + 7 that says: if you know f(n-1), then the next number is found by adding 7 so f(3) = f(2)+7 f(2) is 5, so f(3)= 5+7= 12 so we know f(3)= 12 now find f(4)= f(3)+7 = 12+7= 19 now find f(5) can you do that ?